OpenStudy (anonymous):

Sal currently has an account balance of $2,835.48. He opened the account five years ago with a deposit of $2,310.72. If the interest compounds twice a year, what is the interest rate on the account? Multiple Choice : 2.1% 4.1% 5.9% 8.4%

OpenStudy (johnweldon1993):

Remember that formula from before? \[\large A = P(1 + \frac{ r }{ n })^{nt}\] Well this time..we have A = $2,835.48 P = 2,310.72 N = 2 and T = 5 so we plug everything in..and we get... \[\large 2835.48 = 2310.72(1 + \frac{ r }{ 2 })^{2 \times 5}\] Can you solve for 'r' ...?

OpenStudy (anonymous):

I dont understand the process of the r? can you please explain..

OpenStudy (anonymous):

do i multiply like cross multiply?

OpenStudy (anonymous):

or can i put a number representing r?

OpenStudy (anonymous):

do i plug in each rate and then see if it gives me the answer?

OpenStudy (johnweldon1993):

Sorry about that I was away... Okay well...you CAN do that if you really want to...but a better way would be to rearrange this equation for 'r' First...divide both sides by 2310.72....What do we have after that?

OpenStudy (anonymous):

okay hold on let me divide

OpenStudy (anonymous):

so i divide the left side? now?

OpenStudy (johnweldon1993):

You divide both sides by the 2310.72 \[\large \frac{ 2835.48 }{ 2310.72 } = \frac{ 2310.72(1 + \frac{ r }{ 2 })^{10} }{ 2310.72 }\] like that...

OpenStudy (anonymous):

yes thats what i did and i scratched out the 2310.72 from the right side since it crosses out right?

OpenStudy (johnweldon1993):

Correct...

OpenStudy (anonymous):

now i got 1.227098047 for the left side is that fine?

OpenStudy (johnweldon1993):

Yeah we can keep all those numbers....it'll come out pretty much to the same number at the end... Okay so now we have \[\large 1.227098047 = (1 + \frac{r}{2})^{10} \] Now I think it would be easier to just take the 10th root of both sides...like \(\large \sqrt[10]{n} \) Like that of both sides....so lets do that... \[\large \sqrt[10]{1.227098047} = \sqrt[10]{(1 + \frac{r}{2})^{10}} \] What does this simplify out to...?

OpenStudy (anonymous):

okay let me try this

OpenStudy (anonymous):

left 11.07744577

OpenStudy (anonymous):

do we cross out the 10's in the right side?

OpenStudy (johnweldon1993):

We're going to have to do the left side again but yes the "10's" cancel out on the right...

OpenStudy (johnweldon1993):

\[\huge \sqrt[10]{1.227098047} = 1.0206760553\] Does that look right??

OpenStudy (anonymous):

no ?

OpenStudy (anonymous):

im so sorry im just really confused sorry sorry

OpenStudy (johnweldon1993):

Well remember we are asking ourselves "what number times itself 10 times....= the 1.227098047 This comes out to what I posted....wait are you using a calculator?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

I know exactly what you did...you wrote \[\huge 10\sqrt{1.227098047}\] What this does.....is take the SQUARE ROOT of the number...then multiply it by 10... What that needs to be is... \[\huge \sqrt[10]{1.227098047}\] try typing that in google if you need to..."tenth root of 1.227098047"

OpenStudy (anonymous):

it says its the answer you got before

OpenStudy (anonymous):

1.02067605529

OpenStudy (johnweldon1993):

Right....so that is something you have to watch our for when doing these with your calculator... Okay so now we have: \[\large 1.02067605529 = 1 + \frac{ r }{ 2 }\] Now remember our GOAL is to isolate 'r'...so lets subtract 1 from both sides of the equation... what do we have now?

OpenStudy (anonymous):

.020676055 left r/2 right

OpenStudy (johnweldon1993):

That is correct...and how do we finish solving for 'r'...?

OpenStudy (anonymous):

we divide by 2

OpenStudy (anonymous):

i mean multiply

OpenStudy (johnweldon1993):

Not quite..rememb..... nvm you caught it lol..yes multiply both sides by 2...

OpenStudy (anonymous):

but its not giving the correct answer ...

OpenStudy (johnweldon1993):

You'll see in a second...post what you got...

OpenStudy (anonymous):

0.04135211=r

OpenStudy (johnweldon1993):

Right...now Remember your last question when I said...to convert from a percent to a decimal..you put the number over 100? Well to convert from a decimal to a percent...you MULTIPLY by 100.. so now you get...?

OpenStudy (anonymous):

ohhhhh 4.1

OpenStudy (johnweldon1993):

There you go! :)

OpenStudy (anonymous):

Thank you so much would you mind helping me with one more?

OpenStudy (johnweldon1993):

Oh geez another one >.> haha no just kidding....yeah go ahead and post it :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Ramona deposited $4,190.51 into a savings account with an interest rate of 5.2% compounded twice a year. About how long will it take for the account to be worth $9,000? 33 years, 5 months 29 years, 9 months 14 years, 11 months 20 years, 8 months

OpenStudy (anonymous):

I think its c but im not sure this is a guess

OpenStudy (anonymous):

so can you explain in depth because.. yeah lol

OpenStudy (johnweldon1993):

Lol well okay...first ...Same formula...and lets plug in your values... \[\large 9000 = 4190.51(1 + \frac{ .052 }{ 2 })^{2t}\] Alright...so how did we start off the last problem? what would be step 1?

OpenStudy (anonymous):

divide by 4190.51

OpenStudy (johnweldon1993):

Right...

OpenStudy (anonymous):

2.15 left

OpenStudy (anonymous):

(1+.052/2)^2t

OpenStudy (johnweldon1993):

Okay so... *lets just keep it exact.... 2.1477 left... so \[\large 2.1477 = (1 + \frac{ .052 }{ 2 })^{2t}\] Now do what's inside of the parenthesis..what do you get from that?

OpenStudy (anonymous):

2.1477(1.026)^2t

OpenStudy (johnweldon1993):

Remember the '=' sign inbetween those...but yes.. \[\large 2.1477 = (1.026)^{2t}\] Now...I take it you have gone over "logs" before?

OpenStudy (anonymous):

i honestly dont remember logs

OpenStudy (anonymous):

your probs gunna kill me sorry

OpenStudy (anonymous):

yeah i checked over my work and its 14 and 11 months

OpenStudy (anonymous):

thanks though you were much help :)

OpenStudy (anonymous):

Im so sorry for all teh questions and im dumb slow brain Im really bad at math

OpenStudy (johnweldon1993):

Alright...well I'll show you what to do lol....but remember to go over those :) When you take the natural log (ln) of both sides of this equation...it allows you to bring the exponent containing the variable down in front of the term... aka \[\large 1.026^{2t}\] after taking the natural log we have \[\large 2t \times \ln (1.026)\] And remember we have to do it to the other side as well...so altogher we have \[\large \ln (2.1477) = 2t \times \ln(1.026)\] *just remember to look over those again okay?

OpenStudy (johnweldon1993):

lol I didn't see everything you posted lol...yes it does come out to 14 and 11 months....well good work!